# LASeR: Learning to Adaptively Select Reward Models with Multi-Armed Bandits

### v2

@@ Abstract
Reward Models (RMs) are crucial to aligning large language models (LLMs), but the degree to which an RM specialized to one task (e.g. writing) generalizes to new tasks (e.g. math) is often not known a priori, often making using only one fixed RM to train LLMs suboptimal. However, optimizing LLMs with multiple RMs simultaneously can incur a prohibitively high computational cost and lead to conflicting signals from different RMs that may degrade performance. To address these challenges, we introduce LASeR (Learning to Adaptively Select Rewards), which frames reward model selection as a multi-armed bandit problem, efficiently and iteratively training LLMs using multiple RMs by selecting the most well-suited RM for each instance. On commonsense and math reasoning tasks, we show that LASeR boosts iterative LLM training, improving the absolute average accuracy of Llama-3-8B over three datasets by $2.67\%$ over an ensemble of RM scores while also showing superior efficiency (e.g., a 2×2$\times$2 × speedup). Moreover, on WildChat (open-ended instruction-following tasks), LASeR leads to a $72.69\%$ AlpacaEval win rate over the RM score ensemble baseline. Extending to long-context generation, LASeR improves by $2.96$ F1 points (avg.) on single-document QA tasks and $2.97$ F1 points on few-shot learning over the RM score ensemble baseline with best-of-$n$ sampling. [^1] Code: https://github.com/duykhuongnguyen/LASeR-MAB
11footnotetext: Equal Contribution
@@ 1 Introduction
However, selecting one reward model to guide LLM training can be suboptimal for three main reasons: (1) A single RM may not be generalized to heterogeneous sets of examples. RMs are typically designed to reflect specific objectives and may be trained on offline preference datasets. Thus, an RM that performs well on one dataset or domain may not generalize effectively to others, leading to misaligned outputs across different tasks or domains (Kirk et al., 2023; Chen et al., 2024a; Casper et al., 2023; Gao et al., 2023). For instance, creativity plays a key role in evaluating the quality of a story, whereas correctness is more important in scoring math solutions. (2) RM performance leaderboards (e.g., Lambert et al. (2024)) that rely on human-annotated preferences can have unreliable rankings due to the presence of incorrect and ambiguous preferences (Yu et al., 2024; Hejna et al., 2023). (3) Lastly, over-optimization on one RM can lead to reward hacking problem (Skalse et al., 2022; Rafailov et al., 2024a), even resulting in downstream performance drops.
@@
To mitigate these issues, a prevalent approach is to ensemble multiple reward models (Coste et al., 2023; Eisenstein et al., 2023; Zhang et al., 2024; Ramé et al., 2024). However, these methods also come with significant challenges: as RMs are typically based on LLMs, training with multiple RMs often requires loading and managing several large models simultaneously, which can be computationally expensive, becoming infeasible as models increase in size. Moreover, aggregating multiple RM scores together is susceptible to noisy rewards or conflicting preferences from RMs, especially RMs that are not well-suited for the specific task (Rita et al., 2024). This, in turn, can degrade the quality of the preference data, leading to low-quality updates during training (Wang et al., 2024a). Finally, manually selecting a subset of RMs to combine is a labor-intensive process that involves training many different variants on a combinatorially large set of RM groupings. This underscores the need for more efficient and robust LLM optimization with multiple RMs.
@@
To fill this gap, we introduce Learning to Adaptively Select Rewards (LASeR), that, given a set of RMs, adaptively and efficiently selects a suitable RM for each instance by casting selection as a multi-armed bandit problem (Vermorel and Mohri, 2005; Audibert et al., 2009). Specifically, during training, the RM (arm) is chosen dynamically based on contextual information about the model’s performance and past interactions. The LLM is then fine-tuned based on the RM-annotated data, and the bandit’s parameters are updated accordingly to reflect the performance of the LLM after training on preference data annotated using selected RM (see Fig. 1). By design, LASeR’s adaptive instance-level RM selection (c.f. Sec. 3) addresses the three shortcomings of choosing one reward model: lack of generalization, unreliable rankings, and over-optimization. LASeR eliminates the need for manual tuning or oracle RM selection, with users only tasked with selecting a pool of RMs from a leaderboard like RewardBench (Lambert et al., 2024) without needing to know in advance which RM is best suited for a specific task. Moreover, unlike previous multi-RM methods that do not address conflicting signals from multiple RMs and require simultaneously loading and running multiple RMs (Ramé et al., 2024; Coste et al., 2023) or training an RM for a specific task (Quan, 2024), LASeR selects one RM at a time (Sec. 4), avoiding conflicts and improving efficiency.
@@
Empirically, we demonstrate the effectiveness of LASeR for iteratively training LLMs using multiple RMs on three broad domains: reasoning, instruction-following in text generation, and long-context understanding (Sec. 4.2). We show that on reasoning benchmarks such as StrategyQA (Geva et al., 2021) (testing commonsense reasoning) and GSM8K (Cobbe et al., 2021) (testing math reasoning), LASeR with Llama-3-8B improves absolute accuracy (averaged across 3 datasets) by $1.45\%$ over the best single RM and $2.67\%$ over an ensemble of RM scores. LASeR is also effective on general instruction-following: we show that using LASeR with four strong 7B RMs from RewardBench to finetune Llama-3-8B on a subset of WildChat (Zhao et al., 2024) beats LLMs trained with the best RM and with a baseline that sequentially selects RMs, with $56.34\%$ and $71.45\%$ win rates (respectively) on length-controlled AlpacaEval (Dubois et al., 2024). LASeR also beats the ensemble of RM scores baseline with $72.69\%$ win rates using Llama-3-8B. Moreover, our results show the effectiveness of LASeR’s RM selection strategy at inference-time on long-form generation tasks; on LongBench (Bai et al., 2022), we find LASeR beats the ensemble of RM scores baseline by $2.96$ F1 points on single-document QA tasks and $2.97$ F1 points on few-shot learning when using best-of-$n$ sampling for Llama-3-8B. Our analysis reveals that LASeR is more efficient than sequential multi-RM and RM score ensemble baselines in terms of training time (wall-clock hours) by a factor of 3×3$\times$3 ×, and 2×2$\times$2 ×, respectively, while being more robust to conflicting preferences (Sec. 5).
@@ 2 Related Work
Multiple Reward Ensembles. Recent work explores training LLMs using multiple rewards, often via ensembles (Ramé et al., 2024; Wu et al., 2024; Coste et al., 2023; Zhang et al., 2024; Wang et al., 2024b; Jang et al., 2023; Eisenstein et al., 2023). These methods typically aggregate or align scores across RMs, but suffer from inefficiency (both in terms of training and using multiple RMs) and conflicting rewards (Rita et al., 2024). In contrast, LASeR selects one pretrained RM per training step, avoiding these issues while outperforming ensemble-based baselines (Coste et al., 2023; Wang et al., 2024d). Other approaches train an RM from interpretable objectives using a human-annotated dataset (Wang et al., 2024c) or jointly train multiple task-specific RMs with a sparse router (Quan, 2024). Unlike these methods, LASeR instead uses off-the-shelf RMs to train on LLM-generated outputs, which has shown better performance (Ivison et al., 2024), and dynamically selects RMs from a leaderboard using a multi-armed bandit without any RM training or annotated datasets.
@@
Multi-Armed Bandits (MABs). MABs have been widely applied in ML for optimization and selection tasks (Chen et al., 2013; Li et al., 2010, 2018; Graves et al., 2017), including language models (Pasunuru et al., 2020; Krishnamurthy et al., 2024; Dwaracherla et al., 2024). LASeR uses MABs to dynamically select RMs during training, unlike prior work, which uses MABs to select the annotated samples for a single fixed RM (Dwaracherla et al., 2024) or uses MABs for model selection at test-time without RMs (Nguyen et al., 2024; Li, 2025). This allows LASeR to improve LLMs iteratively without gold labels or explicit optimization for evaluation metrics.
@@ 3 LASeR: Learning to Adaptively Select Rewards
First, we expand on the single-iteration training pipeline with a general reward function (Sec. 3.1). Then, in Sec. 3.2, we describe how LASeR dynamically selects an RM using MAB algorithms, assigning the RM for a given instance or batch. Finally, in Sec. 3.3, we describe cross-iteration training and how we update the MAB parameters. An illustration of LASeR is shown in Fig. 1.
Figure 1: Overview of LASeR. Given the query, the multi-armed bandit selects an RM depending on the underlying query and the bandit’s parameters (based on the usage of each RM and the expected MAB reward). At iteration $m$, the LLM generates multiple responses that are scored based on the selected RM for that query. These responses are ranked into preference pairs, which are then used to fine-tune the model. The same train loss $\mathcal{L}^{m}$ is used to update the parameters of the LLM as well as the MAB for the next iteration, making the entire pipeline iterative.
@@ 3.1 Training LLMs Using a Reward Function
LASeR involves training with multiple RMs using a multi-arm bandit (MAB), which selects one model at a time. Therefore, we first describe how we train LLMs with generated data assuming a single RM; this corresponds to the top-right in Fig. 1 (in blue).
Notation.   Following Yuan et al. (2024b) and Pang et al. (2024), we adopt an iterative training pipeline to finetune the LLM for $M$ iterations. Let $\pi_{m}$ be the LLM at iteration $m$; we assume that we start from an initial pretrained model $\pi_{0}$. Let $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ represent the training inputs, where $x_{i}$ is an input query or prompt. Corresponding to each input query $x_{i}$, we sample a set of $n$ responses from the LLM at the current $m^{\text{th}}$ iteration as $\mathbf{y}_{i}=\{y^{1}_{i},y^{2}_{i},\dots,y^{n}_{i}\}\sim\pi_{m}(y|x_{i})$. Let $R^{\star}\!:\!(y_{i}^{j}|x_{i})\rightarrow\mathbb{R}$ be a reward function that can score an LLM-generated response $y_{i}^{j}$ to a query $x_{i}$ based on how well it aligns with specific task objectives or instructions. Note that $R^{\star}(.)$ can be any reward function and may correspond to a single RM, one of the multiple RMs selected by the MAB (as in our case), or even the true reward.
Generating Preference Pairs.   We evaluate each response $y_{i}^{j}$ using the reward function $R^{\star}(y_{i}^{j}|x_{i})$. By comparing the rewards assigned to different responses, we can form $P$ preference pairs $(y_{i}^{w},y_{i}^{l})$, where $y_{i}^{w}$ is preferred over $y_{i}^{l}$ if $R^{\star}(y_{i}^{w}|x_{i})>R^{\star}(y_{i}^{l}|x_{i})$, thereby building a preference dataset: [^2] Following Pang et al. (2024), we randomly sample $P\!=\!10$ pairs corresponding to each prompt $x_{i}$. For brevity, we omit this in the notation of $\mathcal{D}_{\mathrm{pref}}$; but in our setting $|\mathcal{D}_{\mathrm{pref}}|=P\times|\mathcal{D}|$.
@@
where $r_{i}^{w}=\log\frac{\pi_{m}(y_{i}^{w}\mid x_{i})}{\pi_{m-1}(y_{i}^{w}\mid x_{i})}$, $r_{i}^{l}=\log\frac{\pi_{m}(y_{i}^{l}\mid x_{i})}{\pi_{m-1}(y_{i}^{l}\mid x_{i})}$, $\pi_{m}$ and $\pi_{m-1}$ denotes the LLM in the current iteration $m$ and the previous iteration $m-1$ (used as the reference model in DPO loss). Following Yuan et al. (2024b), we use the standard DPO loss for instruction-finetuning. Following Pang et al. (2024), we use the NLL loss on the preferred responses as an additional regularizer for reasoning tasks, i.e., $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}+\mathcal{L}^{m}_{\textsc{nll}}$. In Appendix C, we show that LASeR outperforms baselines irrespective of the choice of the loss function $\mathcal{L}^{m}$.
@@ 3.2 Bandit Algorithms for Adaptive RM Selection
Sec. 3.1 described the data creation and LLM training procedure for our method when using a general RM (Fig. 1; top-right), which trains the LLM for a single mini-batch. Here, we describe the process by which we adaptively select an RM for each batch of queries using bandit algorithms (shown in Fig. 1-left, in yellow) and update the parameters of the bandit (more details in Sec. A.2).
@@
LASeR uses MABs to dynamically identify the most suitable RM for each query $x_{i}$ and task through exploration while simultaneously training the LLM (since this fine-grained information is not known a priori). Pulling a previously un(der)-explored arm allows the MAB to update its information about the relevance and quality of preference pairs built using that RM via the MAB reward (discussed below). If we over-explore RMs, we might waste time on underperforming RMs, slowing down the overall training progress. On the other hand, if we over-exploit, we might prematurely focus on one RM that seems best initially but is not optimal for all queries or tasks. The exploration-exploitation nature of MAB ensures that LASeR adapts to the specific needs of each query (see Appendix B for further analysis) while improving overall performance.
@@
RM Selection in LASeR.   LASeR uses mini-batch training for each iteration, i.e., we use MABs to select a single RM for a batch of prompts $\mathbf{x}_{m,t}$ for $t^{\text{th}}$ batch or training step of iteration $m$ (total of $T$ steps/batches in each iteration). [^3] Note that LASeR can switch between RMs at the instance level if the batch size is set to 1; however, for the sake of efficiency, we batch instances together both for LASeR and the baselines, as this reduces the computational overhead associated with loading RMs onto the GPU. We provide further discussion in Sec. A.1. Let the set of $K$ reward models (or arms) be denoted by =$\mathcal{R}=\{R_{1},R_{2},\dots,R_{K}\}$, where each $R_{k}$ corresponds to a different RM. We use the negative cumulative train loss function on the batch ($-\mathcal{L}^{m}$) at the given iteration $m$, i.e., $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}+\mathcal{L}^{m}_{\textsc{nll}}$ for reasoning tasks, $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}$ for other task, as mentioned in Sec. 3.1, as the MAB reward. The MAB reward is computed after the model is trained on each batch. Specifically, the MAB reward is the negative training loss (DPO), which depends on how clearly the model learns to prefer the RM’s chosen outputs over the rejected ones. A lower DPO loss or a higher MAB reward corresponds to a larger log-likelihood margin between preferred and dispreferred responses, i.e., $\log\pi_{m}(y_{i}^{w}\mid x_{i})-\log\pi_{m}(y_{i}^{l}\mid x_{i})$ for a query $x_{i}$ and preference pair $(y_{i}^{w},y_{i}^{l})$, indicating that the RM’s feedback helped the model sharpen its rankings and increase confidence in distinguishing between chosen and rejected responses. Thus, an RM that provides more informative and consistent rankings is given a higher reward compared to a less-suited RM providing uninformative rankings.444 Note that since DPO training typically begins from a supervised fine-tuning (SFT) checkpoint, the policy model already has some ability to distinguish good responses from bad ones. In Fig. 5 we show empirically that the model’s margin is typically positive, further supporting this claim.

### v3

@@ Abstract
Reward Models (RMs) are crucial to aligning large language models (LLMs), but the degree to which an RM specialized to one task (e.g. writing) generalizes to new tasks (e.g. math) is often not known a priori, often making using only one fixed RM to train LLMs suboptimal. However, optimizing LLMs with multiple RMs simultaneously can incur a prohibitively high computational cost and lead to conflicting signals from different RMs that may degrade performance. To address these challenges, we introduce LASeR (Learning to Adaptively Select Rewards), which frames reward model selection as a multi-armed bandit problem, efficiently and iteratively training LLMs using multiple RMs by selecting the most well-suited RM for each instance. On commonsense and math reasoning tasks, we show that LASeR boosts iterative LLM training, improving the absolute average accuracy of Llama-3-8B over three datasets by $2.67\%$ over an ensemble of RM scores while also showing superior efficiency (e.g., a $2\times$ speedup). Moreover, on WildChat (open-ended instruction-following tasks), LASeR leads to a $72.69\%$ AlpacaEval win rate over the RM score ensemble baseline. Extending to long-context generation, LASeR improves by $2.96$ F1 points (avg.) on single-document QA tasks and $2.97$ F1 points on few-shot learning over the RM score ensemble baseline with best-of-$n$ sampling. [^1] Code: https://github.com/duykhuongnguyen/LASeR-MAB
@@ 1 Introduction
However, selecting one reward model to guide LLM training can be suboptimal for three main reasons: (1) A single RM may not generalize to heterogeneous sets of examples. RMs are typically designed to reflect specific objectives and may be trained on offline preference datasets. Thus, an RM that performs well on one dataset or domain may not generalize effectively to others, leading to misaligned outputs across different tasks or domains (Kirk et al., 2023; Chen et al., 2024a; Casper et al., 2023; Gao et al., 2023). For instance, creativity plays a key role in evaluating the quality of a story, whereas correctness is more important in scoring math solutions. (2) RM performance leaderboards (e.g., Lambert et al. (2024)) that rely on human-annotated preferences can have unreliable rankings due to the presence of incorrect and ambiguous preferences (Yu et al., 2024; Hejna et al., 2023). (3) Lastly, over-optimization on one RM can lead to reward hacking problem (Skalse et al., 2022; Rafailov et al., 2024a), even resulting in downstream performance drops.
@@
To mitigate these issues, a prevalent approach is to ensemble multiple reward models (Coste et al., 2023; Eisenstein et al., 2023; Zhang et al., 2024; Ramé et al., 2024). However, these methods also come with significant challenges: as RMs are typically based on LLMs, training with multiple RMs often requires loading and managing several large models simultaneously, which can be computationally expensive, becoming infeasible as models increase in size. Moreover, aggregating multiple RM scores together is susceptible to noisy rewards or conflicting preferences from RMs, especially RMs that are not well-suited for the specific task (Rita et al., 2024). This, in turn, can degrade the quality of the preference data, leading to low-quality updates during training (Wang et al., 2024a). Finally, manually selecting a subset of RMs to combine is a labor-intensive process that involves training many different variants on a combinatorially large set of RM groupings. This underscores the need for a more efficient and robust LLM training method with multiple RMs.
@@
To fill this gap, we introduce Learning to Adaptively Select Rewards (LASeR), that, given a set of RMs, adaptively and efficiently selects a suitable RM for each instance by casting selection as a multi-armed bandit problem (Vermorel and Mohri, 2005; Audibert et al., 2009). Specifically, during training, the RM (arm) is chosen dynamically based on contextual information about the model’s performance and past interactions. The LLM is then fine-tuned based on the RM-annotated data, and the bandit’s parameters are updated accordingly to reflect the performance of the LLM after training on preference data annotated using selected RM (see Fig.˜1). By design, LASeR’s adaptive instance-level or batch-level RM selection (c.f. Sec.˜3) addresses the three shortcomings of choosing one reward model: lack of generalization, unreliable rankings, and over-optimization. LASeR eliminates the need for manual tuning or oracle RM selection, with users only tasked with selecting a pool of RMs from a leaderboard like RewardBench (Lambert et al., 2024) without needing to know in advance which RM is best suited for a specific task. Moreover, previous multi-RM methods do not explicitly analyze and address conflicting signals from multiple RMs, and require simultaneously loading and running multiple RMs (Ramé et al., 2024; Coste et al., 2023) or training an RM for a specific task (Quan, 2024). On the other hand, LASeR selects one RM at each training step (Sec.˜4), avoiding conflicts between RMs and improving the overall training efficiency.
@@
Empirically, we demonstrate the effectiveness of LASeR for iteratively training LLMs using multiple RMs on three broad domains: reasoning, instruction-following in text generation, and long-context understanding (Sec.˜4.2). We show that on reasoning benchmarks such as StrategyQA (Geva et al., 2021) (testing commonsense reasoning), GSM8K (Cobbe et al., 2021) (testing math reasoning), and MMLU (Hendrycks et al., 2021b) (testing general knowledge reasoning), LASeR with Llama-3-8B improves absolute accuracy (averaged across 3 datasets) by $1.45\%$ over a baseline that uses best single RM for training and $2.67\%$ over an ensemble of RM scores baseline. LASeR is also effective on general instruction-following: we show that using LASeR with four strong 7B RMs from RewardBench to fine-tune Llama-3-8B on a subset of WildChat (Zhao et al., 2024) beats LLMs trained with the best RM and with a baseline that sequentially selects RMs, with $56.34\%$ and $71.45\%$ win rates (respectively) on length-controlled AlpacaEval (Dubois et al., 2024). LASeR also beats the ensemble of RM scores baseline with $72.69\%$ win rates using Llama-3-8B. Moreover, beyond the use in LLM training, our results demonstrate the effectiveness of LASeR’s RM selection strategy at inference time for long-form generation tasks in reranking LLM responses using multiple RMs; on LongBench (Bai et al., 2022), we find LASeR beats the ensemble of RM scores baseline by $2.96$ F1 points on single-document QA tasks and $2.97$ F1 points on few-shot learning when using best-of-$n$ sampling for Llama-3-8B. Our analysis reveals that LASeR is more efficient than sequential multi-RM and RM score ensemble baselines in terms of training time (wall-clock hours) by a factor of $3\times$, and $2\times$, respectively, while being more robust to conflicting preferences (Sec.˜5).
@@ 2 Related Work
Multiple Reward Ensembles. Recent work explores training LLMs using multiple rewards, often via ensembles (Ramé et al., 2024; Wu et al., 2024; Coste et al., 2023; Zhang et al., 2024; Wang et al., 2024b; Jang et al., 2023; Eisenstein et al., 2023). These methods typically aggregate or align scores across RMs, but suffer from inefficiency (both in terms of training and using multiple RMs) and conflicting rewards (Rita et al., 2024). In contrast, LASeR selects one pretrained RM per training step, avoiding these issues while outperforming ensemble-based baselines (Coste et al., 2023; Wang et al., 2024e). Other approaches train an RM from interpretable objectives using a human-annotated dataset (Wang et al., 2024c) or jointly train multiple task-specific RMs with a sparse router (Quan, 2024). Unlike these methods, LASeR instead uses off-the-shelf RMs to train on LLM-generated outputs, which has shown better performance (Ivison et al., 2024), and dynamically selects RMs from a leaderboard using a multi-armed bandit without any RM training or annotated datasets.
@@
Multi-Armed Bandits (MABs). MABs have been widely applied in ML for optimization and selection tasks (Chen et al., 2013; Li et al., 2010, 2018; Graves et al., 2017), including language models (Pasunuru et al., 2020; Krishnamurthy et al., 2024; Dwaracherla et al., 2024). LASeR uses MABs to dynamically select RMs during training, unlike prior work, which uses MABs to select the annotated samples for a single fixed RM (Dwaracherla et al., 2024) or uses MABs for model selection at test-time without RMs (Nguyen et al., 2024; Li, 2025). This allows LASeR to improve LLMs iteratively without gold labels or explicit optimization for evaluation metrics. To our knowledge, LASeR is the first method to do that in the context of aligning LLMs with multiple RMs.
@@ 3 LASeR: Learning to Adaptively Select Rewards
First, we expand on the single-iteration training pipeline with a general reward function (Sec.˜3.1). Then, in Sec.˜3.2, we describe how LASeR dynamically selects an RM using MAB algorithms, assigning the RM for a given instance or batch. Finally, in Sec.˜3.3, we describe cross-iteration training and how we update the MAB parameters. An illustration of LASeR is shown in Fig.˜1.
Figure 1: Overview of LASeR. Given the query, the multi-armed bandit selects an RM depending on the underlying query and the bandit’s parameters (based on the usage of each RM and the expected MAB reward). At iteration $m$, the LLM generates multiple responses that are scored based on the selected RM for that query. These responses are ranked into preference pairs, which are then used to fine-tune the model. The same training loss $\mathcal{L}^{m}$ is used to update the parameters of the LLM as well as the MAB for the next iteration, making the entire pipeline iterative.
@@ 3.1 Training LLMs Using a Reward Function
LASeR involves training with multiple RMs using a multi-armed bandit (MAB), which selects one model at a time. Therefore, we first describe how we train LLMs with generated data assuming a single RM; this corresponds to the top-right in Fig.˜1 (in blue).
Notation.  Following Yuan et al. (2024b) and Pang et al. (2024), we adopt an iterative training pipeline to fine-tune the LLM for $M$ iterations. Let $\pi_{m}$ be the LLM at iteration $m$; we assume that we start from an initial pretrained model $\pi_{0}$. Let $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ represent the training inputs, where $x_{i}$ is an input query or prompt. Corresponding to each input query $x_{i}$, we sample a set of $n$ responses from the LLM at the current $m^{\text{th}}$ iteration as $\mathbf{y}_{i}=\{y^{1}_{i},y^{2}_{i},\dots,y^{n}_{i}\}\sim\pi_{m}(y|x_{i})$. Let $R^{\star}\!:\!(y_{i}^{j}|x_{i})\rightarrow\mathbb{R}$ be a reward function that can score an LLM-generated response $y_{i}^{j}$ to a query $x_{i}$ based on how well it aligns with specific task objectives or instructions. Note that $R^{\star}(.)$ can be any reward function and may correspond to a single RM, one of the multiple RMs selected by the MAB (as in our case), or even the true reward.
Generating Preference Pairs.  We evaluate each response $y_{i}^{j}$ using the reward function $R^{\star}(y_{i}^{j}|x_{i})$. By comparing the rewards assigned to different responses, we can form $P$ preference pairs $(y_{i}^{w},y_{i}^{l})$, where $y_{i}^{w}$ is preferred over $y_{i}^{l}$ if $R^{\star}(y_{i}^{w}|x_{i})>R^{\star}(y_{i}^{l}|x_{i})$, thereby building a preference dataset: [^2] Following Pang et al. (2024), we randomly sample P=10P\!=\!10 pairs corresponding to each prompt $x_{i}$. For brevity, we omit this in the notation of $\mathcal{D}_{\mathrm{pref}}$; but in our setting $|\mathcal{D}_{\mathrm{pref}}|=P\times|\mathcal{D}|$.
@@
where $r_{i}^{w}=\log\frac{\pi_{m}(y_{i}^{w}\mid x_{i})}{\pi_{m-1}(y_{i}^{w}\mid x_{i})}$, $r_{i}^{l}=\log\frac{\pi_{m}(y_{i}^{l}\mid x_{i})}{\pi_{m-1}(y_{i}^{l}\mid x_{i})}$, $\pi_{m}$ and $\pi_{m-1}$ denote the LLM in the current iteration $m$ and the previous iteration m1m-1 (used as the reference model in DPO loss). Following Yuan et al. (2024b), we use the standard DPO loss for instruction-tuning. Following Pang et al. (2024), we use the NLL loss on the preferred responses as an additional regularizer for reasoning tasks, i.e., $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}+\mathcal{L}^{m}_{\textsc{nll}}$. In Appendix˜C, we show that LASeR outperforms baselines irrespective of the choice of the loss function $\mathcal{L}^{m}$.
@@ 3.2 Bandit Algorithms for Adaptive RM Selection
Sec.˜3.1 described the data creation and LLM training procedure for our method when using a general RM (Fig.˜1; top-right), which trains the LLM for a single mini-batch. Here, we describe the process by which we adaptively select an RM for each batch of queries using bandit algorithms (shown in Fig.˜1-left, in yellow) and update the parameters of the bandit (more details in Sec.˜A.2).
@@
Challenges in Applying MAB to LLM Training.  Although MAB is a promising framework for selecting RMs in LLM training, several challenges remain. First, in typical MAB problems, the arms are often fixed with stable reward distributions. In contrast, during preference fine-tuning of LLMs, the arms are RMs whose outputs interact with and influence the training of the LLM itself. As the LLM is updated, the distribution of responses (and therefore the RM-derived reward signal) shifts. Second, unlike traditional MAB settings where rewards are observed directly, in our setting, no explicit supervision is available to indicate which RM is best per query. Finally, a static MAB setup struggles when new RMs are introduced or when existing RMs are noisy or domain-specific. We describe LASeR in detail, explain how it incorporates MAB into LLM training, and addresses these challenges in the following section.
@@
Exploration and Exploitation of RMs.LASeR uses MABs to dynamically identify the most suitable RM for each query $x_{i}$ and task through exploration while simultaneously training the LLM (since this fine-grained information is not known a priori). Pulling a previously un(der)-explored arm allows the MAB to adaptively update its information about the relevance and quality of preference pairs built using that RM via the MAB reward (discussed below). If we over-explore RMs, we might waste time on underperforming RMs, slowing down the overall training progress. On the other hand, if we over-exploit, we might prematurely focus on one RM that seems best initially but is not optimal for all queries or tasks. The exploration-exploitation nature of MAB ensures that LASeR adapts to the specific needs of each query (see Appendix˜B for further analysis) while improving the overall performance of the model.
@@
RM Selection in LASeR.LASeR uses mini-batch training for each iteration, i.e., we use MABs to select a single RM for a batch of prompts $\mathbf{x}_{m,t}$ for $t^{\text{th}}$ batch or training step of iteration $m$ (total of $T$ steps/batches in each iteration). [^3] Note that LASeR can switch between RMs at the instance level if the batch size is set to 1; however, for the sake of efficiency, we batch instances together both for LASeR and the baselines, as this reduces the computational overhead associated with loading RMs onto the GPU. We provide further discussion in Sec. A.1. Let the set of $K$ reward models (or arms) be denoted by $\mathcal{R}=\{R_{1},R_{2},\dots,R_{K}\}$, where each $R_{k}$ corresponds to a different RM. We use the negative cumulative train loss function on the batch ($\mathcal{L}^{m}$) at the given iteration $m$, i.e., $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}+\mathcal{L}^{m}_{\textsc{nll}}$ for reasoning tasks, $\mathcal{L}^{m}\!=\!\mathcal{L}^{m}_{\textsc{dpo}}$ for other tasks, as mentioned in Sec.˜3.1, as the MAB reward. The MAB reward is computed after the model is trained on each batch. Specifically, the MAB reward is the negative training loss (DPO), which depends on how clearly the model learns to prefer the RM’s chosen outputs over the rejected ones. A lower DPO loss or a higher MAB reward corresponds to a larger log-likelihood margin between preferred and dispreferred responses, i.e., $\log\pi_{m}(y_{i}^{w}\mid x_{i})-\log\pi_{m}(y_{i}^{l}\mid x_{i})$ for a query $x_{i}$ and preference pair $(y_{i}^{w},y_{i}^{l})$, indicating that the RM’s feedback helped the model sharpen its rankings and increase confidence in distinguishing between chosen and rejected responses. Thus, an RM that provides more informative and consistent rankings is given a higher reward compared to a less-suited RM providing uninformative rankings [^4] Note that since DPO training typically begins from a supervised fine-tuning (SFT) checkpoint, the policy model already has some ability to distinguish good responses from bad ones. In Fig. 5 we show empirically that the model’s margin is typically positive, further supporting this claim.. To further validate our MAB reward design, we empirically compare it with alternative learning signals for MAB in Appendix˜C. It is also worth noting that, since the MAB reward is computed based on preference pair data, LASeR can treat any RMs, including learned RM signals or domain-specific metrics as part of its selection framework, thereby enabling generalization and adaptation to different types of RMs (see Appendix˜C for further analysis and experiments).